# CodeRL+: Improving Code Generation via Reinforcement with Execution Semantics Alignment

### v1

@@ Abstract
While Large Language Models (LLMs) excel at code generation by learning from vast code corpora, a fundamental semantic gap remains between their training on textual patterns and the goal of functional correctness, which is governed by formal execution semantics. Reinforcement Learning with Verifiable Rewards (RLVR) approaches attempt to bridge this gap using outcome rewards from executing test cases. However, solely relying on binary pass/fail signals is inefficient for establishing a well-aligned connection between the textual representation of code and its execution semantics, especially for subtle logical errors within the code. In this paper, we propose CodeRL+, a novel approach that integrates execution semantics alignment into the RLVR training pipeline for code generation. CodeRL+ enables the model to infer variable-level execution trajectory, providing a direct learning signal of execution semantics. CodeRL+ can construct execution semantics alignment directly using existing on-policy rollouts and integrates seamlessly with various RL algorithms. Extensive experiments demonstrate that CodeRL+ outperforms post-training baselines (including RLVR and Distillation), achieving a 4.6% average relative improvement in pass@1. CodeRL+ generalizes effectively to other coding tasks, yielding 15.5% and 4.4% higher accuracy on code-reasoning and test-output-generation benchmarks, respectively. CodeRL+ shows strong applicability across diverse RL algorithms and LLMs. Furthermore, probe analyses provide compelling evidence that CodeRL+ strengthens the alignment between code’s textual representations and its underlying execution semantics.00footnotetext: Work done during Xue Jiang and Yihong Dong’s internship at Tongyi Lab.111Our source code will be released at https://github.com/jiangxxxue/CODERLPLUS.
@@ 1 Introduction
In this paper, we propose CodeRL+, which advances standard RLVR training for code generation by incorporating execution semantics alignment. Our approach performs parallel reinforcement learning that jointly optimizes code generation and execution semantics alignment, where the latter repurposes failed exploration programs to analyze their underlying execution semantics by inference how variables propagate during program execution. This integration explicitly aligns the generated code’s textual form with its functional behavior, providing dense learning signals that effectively bridge the gap between textual fluency and execution correctness in LLMs. Crucially, execution semantics alignment employs a on-the-fly training scheme that can be dynamically constructed from code generation rollout programs, requiring no additional data source while evolving with the model’s capabilities.
@@ 3.1 Preliminaries and Definition
where is the importance sampling ratio. is the group-normalized advantage estimate. The KL divergence term prevents the policy from deviating too far from the reference model.
@@ 4.3 Analysis
To investigate the impact of incorporating execution semantics alignment on the model’s internal representations, we conduct a probe experiment. Probes are supervised models trained to predict specific properties from learned representations (Hewitt & Liang, 2019, Lee et al., 2021). In our case, we analyze whether the model’s representations implicitly encode execution semantics, specifically, whether the representations of variables in generated code can predict their runtime values. Our experimental setup (detailed in Appendix B) employs linear regression probes to predict intermediate variable values from the hidden states extracted at variable token positions. We evaluate three models: the base model, GRPO-trained model, and our CodeRL+-trained model. The probe performance, measured by Mean Squared Error (MSE) (Ju et al., 2024) on normalized variable values, serves as a quantitative indicator of how well the model’s internal representations align with actual execution semantics.
@@ Appendix B Setup of Probe Analyses
Appendix D Limitations

### v2

@@ Abstract
While Large Language Models (LLMs) excel at code generation by learning from vast code corpora, a fundamental semantic gap remains between their training on textual patterns and the goal of functional correctness, which is governed by formal execution semantics. Reinforcement Learning with Verifiable Rewards (RLVR) approaches attempt to bridge this gap using outcome rewards from executing test cases. However, solely relying on binary pass/fail signals is inefficient for establishing a well-aligned connection between the textual representation of code and its execution semantics, especially for subtle logical errors within the code. In this paper, we propose CodeRL+, a novel approach that integrates execution semantics alignment into the RLVR training pipeline for code generation. CodeRL+ enables the model to infer variable-level execution trajectory, providing a direct learning signal of execution semantics. CodeRL+ can construct execution semantics alignment directly using existing on-policy rollouts and integrates seamlessly with various RL algorithms. Extensive experiments demonstrate that CodeRL+ outperforms post-training baselines (including RLVR and Distillation), achieving a 4.6% average relative improvement in pass@1. CodeRL+ generalizes effectively to other coding tasks, yielding 15.5% and 4.4% higher accuracy on code-reasoning and test-output-generation benchmarks, respectively. CodeRL+ shows strong applicability across diverse RL algorithms and LLMs. Furthermore, probe analyses provide compelling evidence that CodeRL+ strengthens the alignment between code’s textual representations and its underlying execution semantics.
@@ 1 Introduction
In this paper, we propose CodeRL+, which advances standard RLVR training for code generation by incorporating execution semantics alignment. Our approach performs parallel reinforcement learning that jointly optimizes code generation and execution semantics alignment, where the latter repurposes failed exploration programs to analyze their underlying execution semantics by inferring how variables propagate during program execution. This integration explicitly aligns the generated code’s textual form with its functional behavior, providing dense learning signals that effectively bridge the gap between textual fluency and execution correctness in LLMs. Crucially, execution semantics alignment employs an on-the-fly training scheme that can be dynamically constructed from code generation rollout programs, requiring no additional data source while evolving with the model’s capabilities.
@@ 3.1 Preliminaries and Definition
where is importance sampling ratio, and is the group-normalized advantage estimate.
@@ 4.3 Analysis
To investigate the impact of incorporating execution semantics alignment on the model’s internal representations, we conduct a probe experiment. Probes are supervised models trained to predict specific properties from learned representations (Hewitt and Liang, 2019, Lee et al., 2021). In our case, we analyze whether the model’s representations implicitly encode execution semantics, specifically, whether the representations of variables in generated code can predict their runtime values. Our experimental setup (detailed in Appendix B) employs linear regression probes to predict intermediate variable values from the hidden states extracted at variable token positions. We evaluate three models: the base model, GRPO-trained model, and our CodeRL+-trained model. The probe performance, measured by Mean Squared Error (MSE) (Ju et al., 2024) on normalized variable values, serves as a quantitative indicator of how well the model’s internal representations align with actual execution semantics.
@@ 5 Conclusion
Acknowledgments
This research is supported by the National Natural Science Foundation of China under Grant No. 62192733, 62192730, 62192731, the National Key R&D Program under Grant No. 2023YFB4503801, and the Beijing Major Science and Technology Project under Contract No. Z251100008425005.
@@ Appendix B Setup of Probe Analyses
Appendix D Implementation Details of Execution Semantics Alignment
In implementation, the variable set is restricted to variables of primitive types (e.g., integers, floats, and strings), whose values can be deterministically represented and compared for reward computation. The ground-truth variable states are obtained by executing the program and capturing the final values of target variables upon termination111The variable state extraction can be integrated into the existing execution process used for computing code generation rewards, requiring only a single execution per program and thus introducing minimal overhead.. Furthermore, programs that encounter runtime errors are filtered out, as they do not yield valid execution trajectories. Only programs with semantic errors, i.e., code that executes normally but produces incorrect outputs, are selected for alignment. Such semantic errors represent the majority of code generation errors in LLMs.
Appendix E Additional Benchmark Results
To further evaluate the generalizability of CodeRL+, we conduct additional experiments on two more code generation benchmarks: MBPP (Austin et al., 2021) and LiveBench (White et al., 2025). As shown in Table 4, our method consistently outperforms the base model and GRPO baseline on both datasets in terms of pass@1, demonstrating that CodeRL+ generalizes well beyond the primary evaluation benchmarks.
| Approach | MBPP | LiveBench |
| Qwen2.5-Coder-7B-Instruct | 70.7 | 46.4 |
| GRPO | 72.1 | 46.9 |
| CodeRL+ | 76.2 | 50.6 |
Appendix F Limitations